A bottle opener requires that a force of 35N must be applied to the handle in order to lift the bottle cap 0.90m. the opener has a V.R of 8.0 and an efficiency of 75%. 1. What is the M.A of the opener

2. What force is applied to the bottle cap

3. How far does the handle of the opener move


1
Expert's answer
2021-02-19T11:49:34-0500

Given,

Force required, F=35N

Displacement of handle, "d_l=0.90m"


V.R.=8, "\eta=0.75" ,


(1) Mechanical Advantage, "M.A.=\eta\times V.R."

"=0.75\times8=6"


(2) Forced applied or bottle cap

"F_e=\dfrac{F_l}{MA}"


"=\dfrac{35}{6}=5.83N"


(3) As the bottle cap is in equilibrium so

net torque is zero.

"f_l\times d_l=f_e\times d_e"


"35\times d_l=5.83\times 0.9"


"d_l=\dfrac{5.247}{35}=0.1499\equiv 0.15"


Hence the handle move 0.15m